If Jupiter is 5.2 AU from the Sun, how many kilometers is it from the Sun?
1
Expert's answer
2015-10-05T10:19:16-0400
1 AU=150 000 000 km →5.2 AU=5.2*150*10^6 km=7.8*10^8 km.